# Design of a Quality Management System based on the EU Artificial Intelligence Act

### v1

@@ Abstract
The Artificial Intelligence Act of the European Union mandates that providers and deployers of high-risk AI systems establish a quality management system (QMS). Among other criteria, a QMS shall help to i) identify, analyze, evaluate, and mitigate risks, ii) ensure evidence of compliance with training, validation, and testing data, and iii) verify and document the AI system design and quality. Current research mainly addresses conceptual considerations and framework designs for AI risk assessment and auditing processes. However, it often overlooks practical tools that actively involve and support humans in checking and documenting high-risk or general-purpose AI systems. This paper addresses this gap by proposing requirements derived from legal regulations and a generic design and architecture of a QMS for AI systems verification and documentation. A first version of a prototype QMS is implemented, integrating LLMs as examples of AI systems and focusing on an integrated risk management sub-service. The prototype is evaluated on i) a user story-based qualitative requirements assessment using potential stakeholder scenarios and ii) a technical assessment of the required GPU storage and performance.
@@
EU Artificial Intelligence Act Quality Management System AI Systems Software as a Service Compliance Management
@@ 1 Introduction
Over recent years, the rise of Artificial Intelligence (AI) has introduced rapid advancements and heightened risks, particularly in critical sectors like medicine, finance, and law, where AI increasingly participates in or even controls decision-making processes. To have risks under control, the European Commission introduced its initial draft of the Artificial Intelligence Act (EU AIA) in 2021, which was updated and adopted by the European Parliament on 13 March 2024, accepted by the EU member states on 21 May 2024 and published on 12 July 2024 in the Official Journal of the European Union and will therefore come into legal force on 1 August 2024. It is based on a risk-based approach that categorizes AI systems into distinct risk classes. Specific evidence of safety and security measures must be provided for high-risk AI systems. However, a notable exception applies to general-purpose AI (GPAI) models, as stated by the Future of Life Institute (FLI) [10]. The most prominent examples of GPAIs are large language models (LLMs), which have gained immense popularity since the emergence of OpenAI’s ChatGPT in November 2022. These systems are trained on diverse datasets and applicable across various domains and tasks, making it challenging to neatly classify them into specific risk categories. Consequently, the FLI proposed that GPAIs should provide the same verification and documentation as high-risk AI systems. Even for GPAI models without systematic risk and with less stringent mandatory compliance regulations, implementing the same verification and documentation process as for high-risk AI systems would be beneficial.
Theoretical considerations related to planning and designing risk frameworks by integrating risk standards, human expertise, and audit systems to evaluate AI systems throughout their development lifecycle and during market use have been explored since 2021. The goal is to ensure unified processes and designs as detailed by works such as [7, 24, 27, 28, 33]. Secondly, attention has been directed towards evaluating various characteristics of AI systems, such as performance, explainability, and robustness, and developing techniques to comply with the technical documentation regulations in the EU AIA, as seen in works [1, 3, 11, 26, 31]. However, most approaches focused on metrics to evaluate classification or regression models, often neglecting GPAI models such as LLMs. Furthermore, there is less research on designing and implementing a quality management system (QMS), which is mandatory for high-risk AI systems, as stated in Chapter 3, Section 3, Article 17 EU AIA, and is beneficial for GPAI models. Such a QMS should orchestrate AI development and use in compliance with EU AIA regulations to check and document the AI system’s design, risk, and quality. However, no clear suggestion exists on how such a tool should look. Therefore, the idea is to provide a QMS design as a software as a service (SaaS) web application that assists and involves different stakeholders in the compliance management, verification, and documentation process for AI systems. This QMS should include i) mandatory compliance requirements derived from legal regulations of the EU AIA, ii) technical evaluation metrics to analyze AI systems, and iii) several sub-services, such as a risk management system (Article 9 EU AIA) and a data management and governance system (Article 10 EU AIA). Questions remain about translating legal regulations into software requirements and building a conceptual software architecture model for a QMS. The QMS SaaS should be generic enough to apply to any type of AI system. For that, a generic architecture must be chosen that can easily be extended for different exceptional needs depending on the type of AI system.
@@
This paper aims to fill these gaps by i) constructing legal and structural requirements derived from the EU AIA [9] and developing a conceptual architecture and data model in UML for a QMS based on these requirements. ii) Implementing a first-version prototype QMS that focuses on the integration and adaptation of LLMs as an example of an AI system type. Third, iii) proposing several technical evaluation metrics tailored to LLMs to evaluate performance, explainability, and consistency based on specialized tasks. The paper presents details of the EU AIA in Sect. 2. The method in Sect. 3 outlines the requirements, the QMS architecture, and details of the implemented prototype. The prototype is evaluated and discussed in Sect. 4. The conclusion is provided in Sect. 5.
2 EU Artificial Intelligence Act
@@
The EU Artificial Intelligence Act (EU AIA) [9] is a comprehensive legislative framework regulating the development, deployment, and use of AI systems within the European Union. According to the EU AIA, an AI system “means a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments;”  [^1] Chapter 1 Article 3 para. 1 EU AIA. The EU AIA categorizes AI systems into four different risk classes. Depending on which risk category the AI system is in, specific requirements must be met to be authorized in the EU. The following AI system risk classes exist in the EU AIA:
@@
  • Unacceptable Risk: AI practices pose unacceptable safety, rights, and societal well-being risks. One example would be a social credit system, which evaluates human behavior and actions. These AI systems will be prohibited in the EU.  [^2] Chapter 2 Article 5 EU AIA
@@
  • High Risk: These AI systems are allowed in the EU, but can cause risks to critical areas such as health, safety, and fundamental rights and therefore need to comply with a couple of regulations.  [^3] Chapter 3 Section 1 Article 6 EU AIA
@@
High-risk AI systems must adhere to the most challenging regulations and furnish the EU with documented evidence detailing risk identification, analysis, assessment, mitigation, data management, and data governance. They must be evaluated through their whole lifecycle before entering and post-market. Furthermore, they must demonstrate through technical evaluation that despite being classified as high-risk applications, measures are in place to limit risks effectively. Additionally, GPAI models present a unique challenge in classification under the risk categories. They are versatile and can be employed across various tasks, domains, and purposes, as mentioned by [27]. For that, the FLI came up in 2022 with an article that presents a short list of recommendations that suggested treating GPAI systems the same as high-risk AI systems [10]. They suggested establishing a QMS for GPAIs, the same as for high-risk AI systems, and checking as many regulations as possible from Chapter 3, Section 2. In [28], the authors also plead for categorizing GPAI systems as high-risk AI systems. The EU AIA partly adopted the recommendations from the FLI. It defined a GPAI model as a model “… trained with a large amount of data using self-supervision at scale, that displays significant generality and is capable of competently performing a wide range of distinct tasks regardless of the way the model is placed on the market…”  [^4] Chapter 1 Article 3 para. 63 EU AIA. Based on that definition, GPAI models can also be used in high-risk domains, for high-risk tasks, or integrated into high-risk systems. However, for GPAI models alone, no QMS system as for high-risk AI systems is mandatory; only transparency of used training, validation, and testing data, potential drawbacks, and risk should be documented, which is less strict but could still be changed in future adoptions of the EU AIA. Additionally, the EU AIA defined GPAI models with systematic risk, which is, “… a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole…”  [^5] Chapter 1 Article 3 para. 65 EU AIA. GPAI models with systematic risks are those that use $10^{25}$ FLOPS for training, such as GPT-4, which used $\sim 2*10^{25}$ FLOPS with 1.76 trillion parameters  [^6] Chapter 5 Section 3 Article 55 EU AIA. These models must perform more evaluations and risk assessments and adhere to transparency and documentation obligations. Because most open-source GPAI models are much smaller and trained with less data, they will probably not be treated as having systematic risk.
@@
High-Risk AI Systems Regulations. A Summary of the key regulations for high-risk AI systems  [^7] Chapter 3 Section 2 EU AIA is given. For detailed provisions, refer to the complete articles in the EU AIA [9].
@@
  • Article 9 Risk management system The article states that an RMS should be “established, applied, documented and maintained” (para. 1). The article describes the RMS as an “iterative process” that shall be documented and updated throughout the entire life cycle of the high-risk AI system (para. 2). In addition, the process is specified in the letters. Foreseeable risks shall be identified, analyzed, evaluated, and mitigated (lit. a). Risks that could result in a possible misuse of the AI system shall also be documented and mitigated (lit. b). Risks after placing the system on the market shall be identifiable (lit. c). Targeted risk management measures shall be applied (lit. d). Risk management measures shall be taken in such a way that as few interactions as possible occur (para. 4) but also that residual risks are still considered acceptable (para. 5). To identify risks, the underlying tests shall be carried out under real-world conditions before, during and after the development phase and post-market (para. 6, 7, 8).
@@
  • Article 10 Data and data governance For data used to train high-risk AI systems, it shall be ensured that the training, validation, and testing dataset splits meet the specified quality criteria in para. 2 ff. (para. 1). These datasets shall be governed and managed according to practices suited to the AI system’s intended purpose (para. 2). Key aspects include design choices, data collection and preparation, data suitability, bias examination, and mitigation, and data gaps (para. 2 lit. a - h). Training, validation, and testing datasets shall be relevant, representative, to the best extent, free of errors, and as complete as possible for their intended purpose (para. 3).
@@
  • Article 11 Technical documentation The technical documentation shall be prepared and submitted before the high-risk AI system is placed on the market and shall be continuously updated (para. 1). It shall contain evidence for at least all points listed in Annex IV EU AIA, including all regulations for high-risk AI systems (para. 2).
@@
Article 17 Quality Management System. The QMS aims to structure and plan a process to control at least all listed regulations for high-risk AI systems. The QMS aims to ensure compliance with all EU AIA regulations, verify the AI systems design, and assure its quality. The QMS must be appropriately documented and contain at least the points listed in para. 1 lit. a - m.
GPAI Models Regulations. A Summary of the key regulations for GPAI models  [^9] Chapter 5 Section 2 & Section 3 EU AIA is provided. For detailed provisions, refer to the complete articles in the EU AIA [9]. It is important to note that GPAI models can be included in high-risk AI systems. The obligations for GPAI models and high-risk AI systems must be fulfilled in such cases.
@@
Comments and Reviews. The EU AIA will be legally binding for companies offering AI systems on the European market. However, according to some German legal experts, many uncertainties remain. According to [4], the EU AI Regulation needs more specificity in several areas, making concrete implementation unclear for many requirements. In [29], the author highlights the high and numerous bureaucratic demands for high-risk AI systems. Due to this, compliance checks can require significant human and financial resources [4]. Without innovative approaches to drafting these regulations, only large tech companies will likely have the resources to develop high-risk and generative AI systems. This situation could force small and medium-sized enterprises to leave Europe or avoid developing high-risk AI systems, contradicting the EU AIA Regulation’s objectives and stifling innovation. Additionally, determining if an AI system is classified as high-risk can be very challenging, as described by [8], for example, in the legal tech domain. Another concern is the EU AIA’s lack of alignment and consistency with other legal acts. In [19], it points out that the EU AIA is not fully coordinated with, for example, the Medical Device Regulation, which governs medical devices, including medical software. Additionally, in [5] a German legal expert provides a more comprehensive summary of the EU AIA.
3 Quality Management System for AI Systems
The proposed quality management system (QMS) is designed as a SaaS web application to ensure compliance with EU AIA regulations. It connects directly to an AI system, allowing stakeholders to perform technical quantitative and qualitative checks. The overall idea is to have a single system that verifies the design and quality of the AI system and automatically creates the necessary documentation. The QMS allows for performing technical tests and qualitative assessments from various domain experts. The generated documentation is then sent to national authorities to prove compliance. The first version of the prototype QMS includes two sub-services: a risk management system (RMS) and a data management and governance system (DMDGS). For this prototype implementation, the QMS adapts only LLMs. LLMs are chosen as AI models because of i) their immense popularity, ii) their frequent and domain-independent use, and iii) their open-source availability and easy integration from Hugging Face  [^10] Hugging Face: https://huggingface.co, accessed on 29 July 2024. Even if LLMs are not necessarily high-risk AI systems, in any case, GPAI models, the idea and design of the QMS remain unchanged and can be used for all other types of AI systems. The legal regulations of the EU AIA mostly guide the design and functionality of each sub-service within the QMS. Particular emphasis is placed on the functionality of the RMS sub-service, based on insights from a conducted literature review for designing such a system. In addition to the RMS, the QMS incorporates a DMDGS sub-service. Although the development of the DMDGS was not the primary focus of this first version of prototype QMS, its inclusion should demonstrate that the QMS is designed to integrate multiple different sub-services. The goal is to integrate a separate sub-service for each article for high-risk AI systems. The chapter is structured as follows: It starts with the elicited high-level requirements on how to develop a QMS (cf. Sect. 3.1). Then the architecture and data model of the prototype QMS is described (cf. Sect. 3.2). Lastly, the implementation outcome and the user interface of the QMS and the two sub-services are presented (cf. Sect. 3.3).

### v2

@@ Abstract
The EU AI Act mandates that providers and deployers of high-risk AI systems establish a quality management system (QMS). Among other criteria, a QMS shall help verify and document the AI system design and quality and monitor the proper implementation of all high-risk AI system requirements. Current research rarely explores practical solutions for implementing the EU AI Act. Instead, it tends to focus on theoretical concepts. As a result, more attention must be paid to tools that help humans actively check and document AI systems and orchestrate the implementation of all high-risk AI system requirements. Therefore, this paper introduces a new design concept and prototype for a QMS as a microservice Software as a Service web application. It connects directly to the AI system for verification and documentation and enables the orchestration and integration of various sub-services, which can be individually designed, each tailored to specific high-risk AI system requirements. The first version of the prototype connects to the Phi-3-mini-128k-instruct LLM as an example of an AI system and integrates a risk management system and a data management system. The prototype is evaluated through a qualitative assessment of the implemented requirements, a GPU memory and performance analysis, and an evaluation with IT, AI, and legal experts.
@@
EU AI Act Quality Management System Software as a Service AI Compliance Management
@@ 1 Introduction
Over recent years, the rise of Artificial Intelligence (AI) has introduced rapid advancements and heightened risks, particularly in critical sectors like medicine, finance, and law, where AI increasingly participates in or even controls decision-making processes. To have risks under control, the European Union (EU) Commission introduced the EU AI Act (AIA) in 2021, which came into legal force on 1 August 2024. Since 2021, researchers from the legal tech and information systems community have been working on standardized and unified implementations of the AIA. Many are mainly concerned with developing theoretical ideas, processes, and frameworks on how AI compliance management can be organized in a standardized and organized way (cf. [7, 32, 37, 36]). However, more research is needed on practical tools for efficiently implementing the AIA and involving humans in the AI compliance management process. For instance, a Quality Management System (QMS) responsible for verifying compliance is essential for demonstrating adherence to AIA regulations, particularly for high-risk AI systems. Therefore, this technical report proposes and explains a design concept for an all-in-one QMS solution illustrated in Figure 1.
Figure 1: Proposed design concept for a QMS based on the EU AI Act
The goal is to offer a QMS as a Software as a Service (SaaS) web application, which includes several sub-services that function as independent systems. These sub-services are organized into three core phases of the AI system lifecycle: pre-development, post-development & pre-release, and post-release. Each sub-service can be individually designed to implement specific requirements of the AIA for high-risk AI systems. The QMS shall assist AI system providers in managing compliance, verification, and documentation processes for AI systems. It shall connect with the AI systems to perform technical assessments directly within the QMS. The sub-services shall interact with each other, allowing data to be transferred between them. For example, data from the sub-service addressing Art. 9 AIA can be incorporated into technical documentation, which can then be submitted to national authorities as proof of compliance. Such a QMS can enhance the traceability and reliability of AI compliance management, making it more transparent and allowing for continuous improvement through process mining techniques, as recommended by [39].
@@
This technical report outlines the legal and structural requirements based on the AIA. It presents a conceptual architecture and data model for a QMS aligned with these requirements. It also introduces a first-version prototype QMS, which integrates and adapts a Large Language Model (LLM) as an example of an AI system. The paper presents a summary of the AIA and related work Sect. 2. The method in Sect. 3 outlines the requirements, the QMS architecture, and details of the implemented prototype. The prototype is evaluated and discussed in Sect. 4. The conclusion is provided in Sect. 5.
2 EU Artificial Intelligence Act & Related Work
A summary of the EU AI Act (AIA) is given, followed by related work dealing with the AIA.
2.1 EU Artificial Intelligence Act
@@
The AIA [10] is a comprehensive legislative framework regulating the development, deployment, and use of AI systems within the European Union. According to the AIA, an AI system “means a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments.”  [^1] Chapter 1 Article 3 para. 1 AIA. The AIA categorizes AI systems into four different risk classes. Depending on which risk category the AI system is in, specific requirements must be met to be authorized in the EU. The following AI system risk classes exist in the AIA:
@@
  • Unacceptable Risk: AI practices pose unacceptable safety, rights, and societal well-being risks. One example would be a social credit system, which evaluates human behavior and actions. These AI systems will be prohibited in the EU.  [^2] Chapter 2 Article 5 AIA
@@
  • High Risk: These AI systems are allowed in the EU, but can cause risks to critical areas such as health, safety, and fundamental rights and therefore need to comply with a couple of regulations.  [^3] Chapter 3 Section 1 Article 6 AIA
@@
High-risk AI systems must adhere to the most challenging regulations and furnish the EU with documented evidence detailing risk identification, analysis, assessment, mitigation, data management, and data governance. They must be evaluated through their whole lifecycle before entering and post-market. Furthermore, they must demonstrate through technical evaluation that despite being classified as high-risk applications, measures are in place to limit risks effectively. Additionally, GPAI models present a unique challenge in classification under the risk categories. They are versatile and can be employed across various tasks, domains, and purposes, as mentioned by [35]. For that, the FLI came up in 2022 with an article that presents a short list of recommendations that suggested treating GPAI systems the same as high-risk AI systems [12]. They suggested establishing a QMS for GPAIs, the same as for high-risk AI systems, and checking as many regulations as possible from Chapter 3, Section 2. In [37], the authors also plead for categorizing GPAI systems as high-risk AI systems. The AIA partly adopted the recommendations from the FLI. It defined a GPAI model as a model “… trained with a large amount of data using self-supervision at scale, that displays significant generality and is capable of competently performing a wide range of distinct tasks regardless of the way the model is placed on the market…”  [^4] Chapter 1 Article 3 para. 63 AIA. Based on that definition, GPAI models can also be used in high-risk domains, for high-risk tasks, or integrated into high-risk systems. However, for GPAI models alone, no QMS system as for high-risk AI systems is mandatory; only transparency of used training, validation, and testing data, potential drawbacks, and risk should be documented, which is less strict but could still be changed in future adoptions of the AIA. Additionally, the AIA defined GPAI models with systematic risk, which is, “… a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole…”  [^5] Chapter 1 Article 3 para. 65 AIA. GPAI models with systematic risks are those that use $10^{25}$ FLOPS for training, such as GPT-4, which used $\sim 2*10^{25}$ FLOPS with 1.76 trillion parameters  [^6] Chapter 5 Section 3 Article 55 AIA. These models must perform more evaluations and risk assessments and adhere to transparency and documentation obligations. Because most open-source GPAI models are much smaller and trained with less data, they will probably not be treated as having systematic risk.
@@
High-Risk AI Systems Regulations. A Summary of the key regulations for high-risk AI systems  [^7] Chapter 3 Section 2 AIA is given. For detailed provisions, refer to the complete articles in the AIA [10].
@@
  • Article 9 Risk Management System The article states that a Risk Management System (RMS) should be “established, applied, documented and maintained” (para. 1). The article describes the RMS as an iterative process that shall be documented and updated throughout the entire life cycle of the high-risk AI system (para. 2). In addition, the process is specified in the letters. Foreseeable risks shall be identified, analyzed, evaluated, and mitigated (lit. a). Risks that could result in a possible misuse of the AI system shall also be documented and mitigated (lit. b). Risks after placing the system on the market shall be identifiable (lit. c). Targeted risk management measures shall be applied (lit. d). Risk management measures shall be taken in such a way that as few interactions as possible occur (para. 4) but also that residual risks are still considered acceptable (para. 5). To identify risks, the underlying tests shall be carried out under real-world conditions before, during and after the development phase and post-market (para. 6, 7, 8).
@@
  • Article 10 Data and data governance For data used to train high-risk AI systems, it shall be ensured that the training, validation, and testing dataset splits meet the specified quality criteria in para. 2 ff. (para. 1). These datasets shall be governed and managed according to practices suited to the AI system’s intended purpose (para. 2). Key aspects include design choices, data collection and preparation, data suitability, bias examination and mitigation, and data gaps (para. 2 lit. a - h). Training, validation, and testing datasets shall be relevant, representative, to the best extent, free of errors, and as complete as possible for their intended purpose (para. 3).
@@
  • Article 11 Technical documentation The technical documentation shall be prepared and submitted before the high-risk AI system is placed on the market and shall be continuously updated (para. 1). It shall contain evidence for at least all points listed in Annex IV AIA, including all regulations for high-risk AI systems (para. 2).
@@
Article 17 Quality Management System. The QMS aims to structure and plan a process to control at least all listed regulations for high-risk AI systems. The QMS aims to ensure compliance with all AIA regulations, verify the AI systems design, and assure its quality. The QMS must be appropriately documented and contain at least the points listed in para. 1 lit. a - m.
GPAI Models Regulations. A Summary of the key regulations for GPAI models  [^9] Chapter 5 Section 2 & Section 3 AIA is provided. For detailed provisions, refer to the complete articles in the AIA [10]. It is important to note that GPAI models can be included in high-risk AI systems. The obligations for GPAI models and high-risk AI systems must be fulfilled in such cases.
@@
Comments and Reviews. The AIA is legally binding for companies offering AI systems on the European market. However, according to some German legal experts, many uncertainties remain. According to [5], the EU AI Regulation needs more specificity in several areas, making concrete implementation unclear for many requirements. In [41], the author highlights the high and numerous bureaucratic demands for high-risk AI systems. Due to this, compliance checks can require significant human and financial resources [5]. Without innovative approaches to drafting these regulations, only large tech companies will likely have the resources to develop high-risk and generative AI systems. This situation could force small and medium-sized enterprises to leave Europe or avoid developing high-risk AI systems, contradicting the EU AIA Regulation’s objectives and stifling innovation. Additionally, determining if an AI system is classified as high-risk can be very challenging, as described by [8], for example, in the legal tech domain. Another concern is the AIA’s lack of alignment and consistency with other legal acts. In [26], it points out that the EU AIA is not fully coordinated with, for example, the Medical Device Regulation, which governs medical devices, including medical software.
2.2 Related Work
Literature was searched, published in legal-tech (such as JURIX, AI & Law, ICAIL, and AI and Ethics), and information systems (such as ICPM, and BPM) journals and proceedings to identify theoretical, technical, and practical approaches to efficiently implement the AIA. Approaches for efficient, simplified risk class categorization under the AIA have been presented [18, 15]. Some literature was found describing structured approaches to efficiently implement requirements for high-risk AI systems under the AIA. This includes strategies for Art. 9 AIA [35, 50, 49, 34], as well as methods to enhance AI explainability to make black-box models more transparent, supporting compliance with Art. 13 AIA [44] and Art. 14 AIA [17]. Additionally, work discussing technical metrics to quantitatively assess AI systems’ risks and limitations are presented as required in Art. 9 and 15 AIA [13, 4, 45]. Furthermore, research has been conducted to develop theoretical frameworks and processes for AI auditing, ensuring alignment with legal regulations like the AIA and ethical principles [7, 32, 37, 9, 42]. The business process management community has taken up the AIA in recent research. [36] use process modeling and execution to support reliable AI Fundamental Right Impact Assessment (cf. Art. 29a AIA). [39] emphasize the advantages of process mining for optimizing AI compliance management processes. Furthermore, practical AI compliance management systems like the proposed QMS approach have been introduced. For example, [47] present careAI, a web-based application that provides questions and checklists to evaluate an AI system’s risks, limitations, and benefits. [11] offer capAI, a framework to reliably structure and conduct AI system conformity assessments. In contrast to the design concept proposed in this paper, these tools are more akin to project management frameworks and tools and do not enable technical assessments.
3 Quality Management System Design
The proposed QMS aims to support the AI compliance management process. This section explains the design and implementation in detail. First, the high-level requirements for structuring and implementing the QMS are outlined (cf. Sect. 3.1). Second, the technical evaluation metrics incorporated into the prototype QMS to assess the connected LLM are defined. These metrics include three performance metrics, a gradient-based robustness metric, and a gradient-based explainability metric (cf. Sect. 3.2). Next, the architecture of the QMS, aiming to be modular and scalable, is described, allowing integration with AI systems and incorporating several sub-services (cf. Sect. 3.3). The initial prototype QMS connects to the Phi-3-mini-128k-instruct LLM [1] as an example of an AI system. It includes a risk management sub-service (Art. 9 AIA) and a data management and governance sub-service (Art. 10 AIA). LLMs were chosen for this prototype due to their popularity and the availability of pre-trained, open-source models on Hugging Face [^10] Hugging Face: https://huggingface.co, accessed on 17 October 2024. Lastly, the technical report presents the implementation results and the user interface of the QMS (cf. Sect. 3.4). The prototype’s code can be found in the repository https://github.com/henryk-mustroph/first_version_qmsAIA.git, published on GitHub.